# Motility-Driven Viscoelastic Control of Tissue Morphology in Presomitic Mesoderm

### v1

@@ VI Discussion
Figure 4: Effect of FGF signaling on vertebrate limb formation. (a) Chick embryo (adapted from Ohuchi et al. Ohuchi et al. (1997)): (I) Weak Fgf10 expression in the prospective forelimb mesoderm (arrowheads) is seen around 13 Hamburger and Hamilton (HH) stage. (II) Fgf10 expression in the head region and prospective limb mesoderm (arrowheads) increases at around the 16 $H$ stage. Suggesting Fgf10 plays a key role in limb bud formation. (III) Induction of an ectopic leg-like limb (arrow) following implantation of FGF10-expressing cells in the interlimb region. (b) Zebrafish embryo (adapted from Fischer et al. Fischer et al. (2003)): (I) Wild-type larva at 3 dpf with pectoral fins protruding from the flanks. (II) ika mutant (Fgf24-deficient) larva lacking pectoral fins (asterisks). (c) Mouse embryo (adapted from Min et al.Min et al. (1998)): lateral views of Fgf10+/+, Fgf10+/-, and Fgf10-/- embryos, showing complete limb absence in the knockout.
@@
Our work provides a systematic framework for understanding how cell motility modulates the viscoelastic response of embryonic tissue such as PSM. Using an extended vertex model incorporating active forces, we characterized mechanical behavior under classical rheological protocols such as stress relaxation and oscillatory shear, as well as under spatiotemporally pulsatile perturbations. These studies reveal how motility modulates viscoelastic timescales, which can be used to understand pattern formation in embryonic tissue. We find that motility significantly accelerates stress relaxation, shifting tissue behavior from solid-like to fluid-like by enabling faster junctional rearrangements. In low-motility regimes, residual stress persists over long timescales, leading to elastic storage and global torque under repeated perturbation. In contrast, high-motility tissues reorganize locally, allowing for stress dissipation and morphological adaptation. These findings are consistent with in vivo observations in vertebrate presomitic mesoderm, where posterior regions exhibit fluid-like behavior due to elevated motility, while anterior regions behave more elastically Serwane et al. (2017); Mongera et al. (2018).
A minimal analytical framework captures the essential principles of tissue response under pulsatile forcing. Within a linear viscoelastic description Kelly (2025), the model predicts that long-time deformation decays with increasing perturbation wavenumber ($k_{0}$), reflecting the material’s intrinsic inability to retain high-frequency (cellular-scale) spatial features. This scale-dependent attenuation implies that tissues naturally act as mechanical filters, suppressing fine cellular-scale deformations while preserving longer-wavelength, sub-tissue level patterns.
To test these predictions, we investigated how transient, spatially structured perturbations—motivated by morphogenetic signals—affect long-term tissue morphology in the vertex model. Even when perturbations were applied within the elastic regime ($\mathrm{T_{on}}<\tau_{v}$), tissues accumulated deformation across cycles. The resulting stress led to either alignment or global deformation, depending on the strength of motility and the spatial correlation length of the imposed cue. In agreement with the analytical predictions, sub-tissue–scale perturbations produced persistent morphological patterns in highly motile tissues, whereas cellular-scale cues were effectively dissipated or mechanically filtered out.
@@
Taken together, our results reveal that the spatial correlation of the imposed cue and the level of cellular motility jointly determine the effective bandwidth over which tissues retain and respond to biochemical or mechanical inputs. The intrinsic viscoelasticity of the tissue thus governs how temporal and spatial features of external stimuli are integrated during morphogenesis. Notably, this framework establishes a mechanistic link between molecular signaling and large-scale mechanics: pathways such as FGF can modulate cell motility to tune viscoelastic responses, thereby shaping collective cell dynamics and morphogenetic outcomes.
FGF signaling has been proven to be a key controller of vertebrate limb development. In chick embryos, Fgf10 expression emerges early in the prospective forelimb mesoderm, initially weak around stage 13 (Hamburger and Hamilton (HH) 13) and becoming stronger by stage 16 (HH16), coinciding with the onset of limb bud outgrowth Ohuchi et al. (1997). Functional perturbation experiments demonstrate causality: implantation of FGF10-expressing cells in the interlimb region induces ectopic, leg-like structures, underscoring FGF10’s sufficiency in initiating bud formation Ohuchi et al. (1997). This role is conserved in zebrafish through Fgf24, where wild-type larvae develop prominent pectoral fins. At the same time, the ika mutant (lacking Fgf24) completely fails to initiate fin outgrowth Fischer et al. (2003), highlighting the requirement of FGF signaling for epithelial and mesenchymal motility. Similarly, in mouse embryos wild-type (Fgf10+/+) and heterozygotes (Fgf10+/-) develop normal limbs, homozygous knockouts (Fgf10-/-) show complete limb absence Min et al. (1998). Our results suggest that this FGF-mediated limb formation across species is governed by a localised increase in motility in otherwise static regions (e.g., lateral plate), which modulates the tissue’s viscoelastic properties locally, enabling directed outgrowth and limb formation.
Moreover, during embryogenesis, tissues are exposed to pulsatile mechanical cues spanning a wide range of spatial and temporal scales. Internally, many morphogenetic processes are driven by periodic contractile activity, for example, actomyosin pulses that drive apical constrictions Sutherland and Lesko (2020) and the pulsed forces that underlie dorsal closure Solon et al. (2009), as well as oscillatory gene expression that governs boundary formation during somitogenesis Hubaud and Pourquié (2014).
Tissues also experience rhythmic external forces. In early vertebrate embryos, the developing heart begins rhythmic contractions even before circulation, sending mechanical pulses across millimeter scales that can influence adjacent tissues, including the PSM Hornberger and Sahn (2007); Watanabe-Asaka et al. (2014); Tyser and Srinivas (2020). Similarly, cilia-driven nodal flows generate oscillatory shear forces that not only direct left–right symmetry breaking but may also mechanically stimulate neighboring mesodermal regions Shinohara and Hamada (2017). Together, these internally generated and externally applied rhythmic inputs shape local cell behaviors and coordinate tissue-scale morphogenetic events such as axis elongation, boundary formation, and lumen expansion.
@@
Also, the pulsatile force analysis developed here broadly applies beyond developmental morphogenesis. Many epithelial tissues in physiological settings experience rhythmic or cyclic mechanical loading. For example, the alveolar epithelium in the lungs undergoes continuous stretching and relaxation during normal breathing and mechanical ventilation, where the frequency and amplitude of deformation vary with respiration patterns Trepat et al. (2004); Tschumperlin et al. (2000); Geiger et al. (2009). The intestinal epithelium is rhythmically deformed by peristaltic contractions that generate traveling waves of compression and extension along the gut Wang et al. (2018). Cardiac tissues are subjected to high-frequency cyclic strain during heartbeat cycles, with the heart epithelium enduring over 2.5 billion contraction and expansion events over an average human lifetime Papafilippou et al. (2025). Other examples include bladder epithelial stretch during filling and voiding cycles, and uterine epithelium undergoing deformation during contractions.
In contexts such as mechanical ventilation, where the alveolar epithelium experiences cyclic stretching, deformation buildup over a long period may contribute to ventilator-induced lung injury. Conversely, tuning the frequency and amplitude of mechanical input to match tissue relaxation dynamics or modulating tissue fluidity through physiological or pharmacological interventions may reduce stress accumulation and preserve integrity. Similarly, our model predicts that tissues subject to rhythmic deformation selectively accumulate or dissipate stress depending on the match between input parameters and motility-driven remodeling capacity. This mechanistic framework thus links cyclic input characteristics with emergent morphological and stress responses, offering insight into how tissues can be tuned for robustness under repetitive mechanical loading.
@@
This framework provides a versatile platform for investigating time-dependent pattern formation and stress encoding in active viscoelastic media. It can be readily extended to incorporate features such as anisotropic tension generation, active nematic alignment, and mechano-chemical feedback, thereby enabling the exploration of more complex viscoelastic behaviors relevant to morphogenesis, organoid mechanics, and the design of programmable synthetic tissues.
@@ VII Acknowledgements
These simulations were performed on the Param Seva supercomputers through the National Supercomputing Mission and the IITH Kanad clusters. S.I. acknowledges the Prime Minister Research Fellowship (PMRF, ID-2002732) for financial support through a research fellowship. M.S.R. acknowledges SERB(India) for financial support. A. G. acknowledges SERB-DST (India) Projects MTR/2022/000232, CRG/2023/007056-G, DST (India) grant no. DST/NSM/R&D_HPC_Applications/2021/05 and grant no. SR/FST/PSI-215/2016, and IITH for Seed Grant No. IITH/2020/09 for financial support.
Supplementary Material
S1 Materials and Methods
S1.1 Vertex Model
The energy functional and force for vertex model is given by,
@@
The total force due to this energy functional on a vertex is given as:
$\mathbf{F}=\begin{bmatrix}F_{x}\\ F_{y}\end{bmatrix}=-\nabla_{v}(U_{c})=-2\lambda(A_{c}-A_{0})\begin{bmatrix}\frac{\partial Ac}{\partial x_{v}}\\ \frac{\partial Ac}{\partial y_{v}}\end{bmatrix}$ (S2)
@@
$A_{c}=\frac{1}{2}\sum_{v=1}^{N}\left(x_{v}y_{v+1}-y_{v}x_{v+1}\right),$
where $v+1$ is the next vertex index, and the indices are cyclic (i.e., $x_{N+1}=x_{1}$).
The gradient of $A_{c}$ with respect to $x_{v}$ is:
$\frac{\partial A_{c}}{\partial x_{v}}=\frac{1}{2}\left(y_{v+1}-y_{v-1}\right),$
where $v+1$ and v1v-1 refer to the next and previous vertices, respectively.
Similarly, the gradient of $A_{c}$ with respect to $y_{v}$ is:
$\frac{\partial A_{c}}{\partial y_{v}}=\frac{1}{2}\left(x_{v-1}-x_{v+1}\right).$
The perimeter of a polygon is:
$P_{c}=\sum_{v=1}^{N_{v}}\sqrt{\left(x_{v+1}-x_{v}\right)^{2}+\left(y_{v+1}-y_{v}\right)^{2}}.$
The gradient of $P_{c}$ with respect to $x_{v}$ involves contributions from both adjacent edges ($v-1\to v$ and $v\to v+1$):
$\frac{\partial P}{\partial x_{v}}=\frac{x_{v}-x_{v-1}}{\sqrt{\left(x_{v}-x_{v-1}\right)^{2}+\left(y_{v}-y_{v-1}\right)^{2}}}$
$+\frac{x_{v}-x_{v+1}}{\sqrt{\left(x_{v+1}-x_{v}\right)^{2}+\left(y_{v+1}-y_{v}\right)^{2}}}$
Similarly, the gradient of $P_{c}$ with respect to $y_{v}$ is:
$\frac{\partial P_{c}}{\partial y_{v}}=\frac{y_{v}-y_{v-1}}{\sqrt{\left(x_{v}-x_{v-1}\right)^{2}+\left(y_{v}-y_{v-1}\right)^{2}}}$
$+\frac{y_{v}-y_{v+1}}{\sqrt{\left(x_{v+1}-x_{v}\right)^{2}+\left(y_{v+1}-y_{v}\right)^{2}}}$
@@
$\frac{\partial A_{c}}{\partial\mathbf{r}}=\begin{bmatrix}\frac{\partial A_{c}}{\partial x_{v}}\\ \frac{\partial A_{c}}{\partial y_{v}}\end{bmatrix}=\frac{1}{2}\begin{bmatrix}y_{v+1}-y_{v-1}\\ x_{v-1}-x_{v+1}\end{bmatrix}$ (S3)
$\frac{\partial P_{c}}{\partial\mathbf{r}}=\begin{bmatrix}\frac{\partial P_{c}}{\partial x_{v}}\\ \frac{\partial P_{c}}{\partial y_{v}}\end{bmatrix}=\begin{bmatrix}\frac{x_{v}-x_{v-1}}{\sqrt{\Delta x_{v-1}^{2}+\Delta y_{v-1}^{2}}}+\frac{x_{v}-x_{v+1}}{\sqrt{\Delta x_{v+1}^{2}+\Delta y_{v+1}^{2}}}\\ \frac{y_{v}-y_{v-1}}{\sqrt{\Delta x_{v-1}^{2}+\Delta y_{v-1}^{2}}}+\frac{y_{v}-y_{v+1}}{\sqrt{\Delta x_{v+1}^{2}+\Delta y_{v+1}^{2}}}\end{bmatrix}$ (S4)
@@
To make the cells motile we have added random white noise force $\xi_{c,v}(t)$ with $\langle\xi_{c,v}(t)\rangle=0$ and $\langle\xi_{c,v}(t)\xi_{c^{\prime},v^{\prime}}(t^{\prime})\rangle=2\mathcal{M}\delta_{cc^{\prime}}\delta_{vv}\delta(t-t^{\prime})$) to each vertex position($v$ of a cell $c$) that gives rice to random motion to the cells.
$\eta{\bf\dot{r}}^{v}_{c}=-{\bf\nabla}_{v}(U_{c})+{\boldsymbol{\xi}}_{c,v}(t)$ (S5)
This model incorporates cell junctional rearrangements (T1 transitions) and allows cells to detach from the tissue through T2 transitions or through multiple T1 transitions.
Supplementary Figure S1: (a) A representative epithelial tissue configuration obtained from vertex model simulations. (b) Schematic illustration of a T1 transition involving neighbor exchange. (c) Schematic of a T2 transition representing cell extrusion.
The parameters involved in equation S5 was non dimensionalized using the length scale $L\sim\sqrt{A_{0}}$ and time scale $\frac{1}{\sqrt{\lambda A_{0}}}$. Both the values of $\lambda$ and $A_{0}$ were kept $1$ throughout all the simulations. The deterministic part of equation S5 was solved by implicit Euler time integration method and the stochastic part was solved by Weiner method with time step, $dt=2.5\times 10^{-3}$.
@@
We have calulated tissue circularity by finding out the convex hull of the cell centres of the final structure. Then calculating the area ($A$) and perimeter($P$) of the given QHull the circularity is defined as,
@@
Supplementary Figure S2: Using convex hull to evaluate the circularity of tissue morphology.
@@
Where , $W(x)$ is a Heaviside step function given by, $W(x<0)=1$. It shows the rate of radial movement of a cell from its initial position. The relaxation time for a tissue is defined by the time it takes for $Q(t)$ to become $\frac{1}{e}^{th}$ of its initial value.
@@
Supplementary Figure S3: Cumulative count of T1 transitions ocer time. Higher motility promotes T1 transitions.
@@
Where $\Pi_{c}=-\frac{\partial U_{c}}{\partial A_{c}}$ is the hydrostatic pressure and $\boldsymbol{T}_{e}=\frac{\partial U_{c}}{\partial\boldsymbol{l}_{e}}$ is the line tension/ shear stress term. This leads to the expression,
@@
$\sum_{v=1}^{N_{v}}\begin{pmatrix}\Delta x^{2}_{v+1}&\Delta x_{v+1}\Delta y_{v+1}\\ \Delta y_{v+1}\Delta x_{v+1}&\Delta y^{2}_{v+1}\end{pmatrix}\frac{1}{\sqrt{\Delta x^{2}_{v+1}+\Delta y^{2}_{v+1}}}$
@@
A stress relaxation experiment involves applying an external deformation to the system and observing how the internal stress dissipates over time. To achieve this in the simulation, we start form a random voronoi configuration and let the system relax just by it’s own properties governed by the vertex Hamiltonian, and motility of the cells. Then we apply an affine shear strain ${\bf\gamma}(t)=\left(\begin{array}[]{cc}1&\epsilon\\ 0&1\end{array}\right)$ to the system and maintain the shear strain by keeping the boundary of the tissue fixed at a particular level of strain. We examine the relaxation of the shear stress of the bulk tissue, far away from the boundary.
@@
We have applied oscillatory shear strain to the system with varying frequency of the oscillation and measured the stress response of the system. From this we calculated Storage($G^{\prime}$) and Loss modulus($G^{\prime\prime}$) given by the respectively in-phase and out-of-phase response of stress to the applied strain.
@@
Thus the we have to find out the Fourier coefficients corresponding to the input frequency $\omega_{0}$ from the Fourier series of $\sigma(t)$. One way is to do a Fast Fourier Transform (FFT) numerically and calculate the coefficients. But this requires large sampling data and can be errornous sometimes. To avoid that we have used an alternate calculation follows :
@@
Let’s say $n=n_{0}$ corresponds to the input frequency. So we have to find the coefficients of $\cos(\omega_{n_{0}}t)$ and $\sin(\omega_{n_{0}}t)$ i.e $a_{n_{0}}$ and $b_{n_{0}}$.
Now, if we integrate equation S13 to a time $T$ or multiply it with $\cos(\omega_{n_{0}}t)$ or $\sin(\omega_{n_{0}}t)$ and then integrate, we get three equations involving $a_{0}$, $a_{n_{0}}$ and $b_{n_{0}}$ as three unknowns which we can solve as system of linear equations. These equations looks like,
@@
In equation S14, the other terms where $n\neq n_{0}$ has been ignored as they would not contribute much. In equation S14, left hand side integrals are calculated using numerical integration scheme. Thus, we can calculate $a_{0}$, $a_{n_{0}}$ and $b_{n_{0}}$, and from that we can get $G^{\prime}=\frac{b_{n_{0}}}{\epsilon_{0}}$ and $G^{\prime\prime}=\frac{a_{n_{0}}}{\epsilon_{0}}$.

### v2

@@ VI Discussion
Figure 4: Effect of FGF signaling on vertebrate limb formation. (a) Chick embryo (adapted with permission from Ohuchi et al. ohuchiMesenchymalFactorFGF101997): (I) Weak Fgf10 expression in the prospective forelimb mesoderm (arrowheads) is seen around 13 Hamburger and Hamilton ($H$) stage. (II) Fgf10 expression in the head region and prospective limb mesoderm (arrowheads) increases at around the 16 $H$ stage. Suggesting Fgf10 plays a key role in limb bud formation. (III) Induction of an ectopic leg-like limb (arrow) following implantation of FGF10-expressing cells in the interlimb region. (b) Zebrafish embryo (adapted with permission from Fischer et al. fischerZebrafishFgf24Mutant2003): (I) Wild-type larva at 3 dpf with pectoral fins protruding from the flanks. (II) ika mutant (Fgf24-deficient) larva lacking pectoral fins (asterisks). (c) Mouse embryo (adapted with permission from Min et al.minFgf10RequiredBoth1998): lateral views of Fgf10+/+, Fgf10+/-, and Fgf10-/- embryos, showing complete limb absence in the knockout. (d) Simulations using active vertex model tissue to examine the role of FGF-induced motility in limb bud initiation. Starting from an initially compressed configuration, four localized motility hotspots (I) trigger protrusive outgrowths (II), which grow further resembling early limb bud formation (III). The colorbar represents motility value normalised with the maximum used for the posterior PSM.
@@
Our work provides a unified framework for understanding how cell motility modulates the viscoelastic response of embryonic tissues such as the presomitic mesoderm. Using an active vertex model calibrated to PSM explant relaxation measurements, we reproduced the motility-dependent mechanical behavior of PSM explants and uncovered how active fluctuations reshape tissue-level rheology. By characterizing the model under classical rheological protocols, such as stress relaxation and oscillatory shear, we identified the intrinsic mechanical timescales that govern morphogenetic patterning. These analyses show that motility accelerates stress relaxation by enhancing junctional rearrangements, effectively shifting the tissue from solid-like to fluid-like behavior. In low-motility regimes, where rearrangements are rare, residual stresses persist, and the tissue responds elastically over long timescales. In contrast, highly motile tissues rapidly reorganize to dissipate stress and adapt their morphology. This motility-dependent viscoelastic transition is consistent with in vivo observations in vertebrate PSM, where the posterior domain exhibits fluid-like behavior due to elevated cellular motility, while the anterior region remains more elastic serwaneVivoQuantificationSpatially2017a; mongeraFluidtosolidJammingTransition2018.
A fallout of this motility-driven viscoelasticity is in embryonic tissues, which encounter rhythmic mechanical cues across a range of spatiotemporal scales, including periodic actomyosin contractions sutherland2020pulsed; solon2009pulsed, oscillatory gene-expression rhythms during somitogenesis hubaud2014signalling, early heartbeat-driven deformations hornberger2007rhythm; watanabe2014regular; tyser2020first, and cilia-generated oscillatory flows shinohara2017cilia. These cues have diverse origins but share a common physical role: they impose time-varying stresses that must be integrated by a viscoelastic, actively remodeling material.
To understand how such rhythmic cues sculpt tissue architecture, we developed a linear viscoelasticity-based analytical framework kellyMechanicsLectureNotes2025. The model predicts that the tissue’s long-time deformation decreases sharply with increasing perturbation wavenumber $k_{0}$, reflecting a fundamental physical constraint: tissues cannot retain fine, cell-scale spatial features because these high–wavenumber inputs are rapidly dissipated through local junctional rearrangements. In effect, the tissue acts as a mechanical filter that selectively preserves long-wavelength, sub–tissue–scale cues while suppressing short-wavelength fluctuations. Guided by this principle, we examined the response of the active vertex tissue to pulsatile, spatially structured forcing. Even when each pulse was applied within the nominally elastic regime ($\tau_{\mathrm{on}}\ll\tau_{V}$), repeated cycles accumulated deformation in a motility-dependent manner—driving coherent alignment and pattern formation for long-wavelength inputs, but producing only transient, local distortions for short-wavelength perturbations. Consistent with the analytical predictions, highly motile tissues generated robust, persistent morphological patterns in response to long-wavelength cues, whereas cellular-scale perturbations were rapidly dissipated, leaving no lasting imprint.
@@
Taken together, our results indicate that the spatial structure of an imposed cue and the local level of cellular motility jointly determine the bandwidth over which tissues retain and respond to biochemical or mechanical inputs. The intrinsic viscoelasticity of the tissue, therefore, governs how temporal and spatial features of external stimuli are integrated during morphogenesis. This framework provides a mechanistic bridge between molecular signaling and tissue mechanics: pathways such as FGF can modulate motility to tune viscoelastic responses, thereby shaping collective cell behaviors and morphogenetic flows.
One of the biological contexts where this connection between tissue rheology and external mechanical cues can play a major role is in the FGF signaling-driven initiation of the vertebrate limb bud. In chick embryos, Fgf10 expression emerges in the presumptive forelimb mesoderm between HH13 and HH16, coinciding with the onset of limb-bud outgrowth ohuchiMesenchymalFactorFGF101997 (Fig. 4(a) I,II). Experimental overexpression studies suggest that implanting Fgf10-producing cells in the interlimb region is sufficient to induce ectopic limb-like structures ohuchiMesenchymalFactorFGF101997 (Fig. 4(a) III). This role is conserved across vertebrates: zebrafish fgf24 mutants fail to initiate pectoral fin outgrowth fischerZebrafishFgf24Mutant2003 (Fig. 4(b)), and Fgf10-/- mice lack limb buds entirely minFgf10RequiredBoth1998 (Fig. 4(c)). Collectively, these studies demonstrate that localized FGF activity provides a spatially confined signal that initiates bud formation.
To examine how such molecular cues could physically generate tissue-scale deformation, we used an active vertex model to simulate a mildly compressed epithelial sheet and introduced four spatially localized motility hotspots representing regions of elevated FGF activity. Under constrained extension along one axis, these hotspots produced persistent lateral cell outflow and drove a protrusive deformation reminiscent of early limb-bud emergence (Fig. 4(d)). This minimal simulation illustrates how localized, FGF-driven increases in motility can convert molecular cues into emergent mechanical behaviors, offering a mechanistic interpretation that complements experimental observations across vertebrate systems.
@@
Finally, the pulsatile-force analysis developed here has implications well beyond developmental morphogenesis. Many epithelial tissues in physiological settings experience rhythmic or cyclic mechanical loading. The alveolar epithelium, for example, undergoes continuous stretch–relax cycles during breathing and mechanical ventilation trepat2004viscoelasticity; tschumperlin2000deformation; geiger2009tubulin. The intestinal epithelium is periodically deformed by peristaltic waves wang2018bioengineered, and cardiac epithelia sustain high-frequency cyclic strain over billions of heartbeat cycles PAPAFILIPPOU2025102511. Similar rhythmic deformations occur in the bladder, uterus, and other mechanically active organs. These contexts underscore that pulsatile inputs are not exceptional but represent a ubiquitous mode of mechanical forcing in living tissues.
Our framework shows that whether such cyclic cues accumulate into lasting morphological changes or are efficiently dissipated depends on the interplay between input frequency and the tissue’s motility-controlled relaxation dynamics. In settings such as mechanical ventilation, for instance, mismatches between driving frequency and intrinsic relaxation times may promote long-time deformation buildup, contributing to ventilator-induced injury. Conversely, adjusting deformation frequency or enhancing tissue fluidity—whether physiologically or through intervention—may reduce stress accumulation and preserve epithelial integrity. More broadly, our results indicate that tissues act as frequency-dependent mechanical filters, selectively retaining or erasing cyclic perturbations based on their viscoelastic state.
@@
Overall, this work provides a general mechanistic framework for understanding how active, viscoelastic tissues integrate spatially and temporally structured mechanical cues. Beyond explaining observed behaviors in the PSM and limb initiation, the model offers a versatile platform for probing time-dependent pattern formation, stress encoding, and mechanical robustness in active biological materials. Extensions incorporating anisotropic tension generation, active nematic alignment, or mechano-chemical feedback will enable exploration of richer behaviors relevant to morphogenesis, organoid mechanics, and the engineering of programmable synthetic tissues.
@@ VII Acknowledgements
These simulations were performed on the Paramseva supercomputers through the National Supercomputing Mission and the IITH Kanad clusters. S.I. acknowledges the Prime Minister Research Fellowship (PMRF, ID-2002732) for financial support through a research fellowship. M.S.R. acknowledges SERB (India) project SRG/2021/001020 and IIT Hyderabad for financial support. A. G. acknowledges SERB-DST (India) Projects MTR/2022/000232, CRG/2023/007056-G, DST (India) grant no. DST/NSM/R&D_HPC_Applications/2021/05 and grant no. SR/FST/PSI-215/2016, and IITH for Seed Grant No. IITH/2020/09.
Motility-Driven Viscoelastic Control of Tissue Morphology in Presomitic Mesoderm
(Supplementary Information)
Sahil Islam1, Mohd. Suhail Rizvi2, Anupam Gupta1
1Department of Physics, Indian Institute of Technology Hyderabad, India
2Department of Biomedical Engineering, Indian Institute of Technology Hyderabad, India
S1 Methods
S1.1 Active Vertex Model
The energy functional for a single cell ($c$) in the vertex model is given by,
@@
Where $A_{c}$ is the area and $P_{c}$ is the perimeter of a cell (Fig. S1(a)) and $A_{0}$ is the target area of a cell.
The total force on a vertex ($v$) of a cell ($c$) is given as:
$\mathbf{F}=-\nabla_{v}(U_{c})=\begin{bmatrix}F_{x}\\ F_{y}\end{bmatrix}=-2\lambda(A_{c}-A_{0})\begin{bmatrix}\frac{\partial Ac}{\partial x_{v}}\\ \frac{\partial Ac}{\partial y_{v}}\end{bmatrix}$ (S2)
@@
$A_{c}=\frac{1}{2}\sum_{v=1}^{N_{c}}\left(x_{v}y_{v+1}-y_{v}x_{v+1}\right),$
where $N_{c}$ is the number of vertex in the cell $c$ and $v+1$ is the next vertex index, and the indices are cyclic (i.e., $x_{N_{c}+1}=x_{1}$ ).
The perimeter of a polygon is given by:
$P_{c}=\sum_{v=1}^{N_{c}}\sqrt{\left(x_{v+1}-x_{v}\right)^{2}+\left(y_{v+1}-y_{v}\right)^{2}}.$
@@
$\frac{\partial A_{c}}{\partial\mathbf{r}_{v}}=\begin{bmatrix}\frac{\partial A_{c}}{\partial x_{v}}\\ \frac{\partial A_{c}}{\partial y_{v}}\end{bmatrix}=\frac{1}{2}\begin{bmatrix}y_{v+1}-y_{v-1}\\ x_{v-1}-x_{v+1}\end{bmatrix}$ (S3)
$\frac{\partial P_{c}}{\partial\mathbf{r}_{v}}=\begin{bmatrix}\frac{\partial P_{c}}{\partial x_{v}}\\ \frac{\partial P_{c}}{\partial y_{v}}\end{bmatrix}=\begin{bmatrix}\frac{x_{v}-x_{v-1}}{\sqrt{\Delta x_{v-1}^{2}+\Delta y_{v-1}^{2}}}+\frac{x_{v}-x_{v+1}}{\sqrt{\Delta x_{v+1}^{2}+\Delta y_{v+1}^{2}}}\\ \frac{y_{v}-y_{v-1}}{\sqrt{\Delta x_{v-1}^{2}+\Delta y_{v-1}^{2}}}+\frac{y_{v}-y_{v+1}}{\sqrt{\Delta x_{v+1}^{2}+\Delta y_{v+1}^{2}}}\end{bmatrix}$ (S4)
@@
To make the cells motile we have added force as random white noise $\xi_{v,\alpha}(t)$ with $\langle\xi_{v,\alpha}(t)\rangle=0$ and $\langle\xi_{v,\alpha}(t)\xi_{v^{\prime},\alpha^{\prime}}(t^{\prime})\rangle=2\mathcal{M}(\mathbf{r})\eta\delta_{vv^{\prime}}\delta_{\alpha\alpha^{\prime}}\delta(t-t^{\prime})$ ) to each vertex position($v$) and $\alpha$ represents the direction of the force.
$\eta\dot{\mathbf{r}}_{v}=-\nabla_{v}U+\boldsymbol{\xi}_{v}(t).$ (S5)
This model incorporates cell junctional rearrangements (T1 transitions) and allows cells to detach from the tissue through T2 transitions (Fig. S1(b),(c)).
Figure S1: (a) A representative epithelial tissue configuration obtained from vertex model simulations. (b) Schematic illustration of a T1 transition involving neighbor exchange. (c) Schematic of a T2 transition representing cell extrusion.
The parameters involved in equation S5 were non-dimensionalized using the length scale $L\sim\sqrt{A_{0}}$ and time scale $\frac{1}{\sqrt{\lambda A_{0}}}$. Both the values of $\lambda$ and $A_{0}$ were kept $1$ throughout all the simulations. The deterministic part of equation S5 was solved by the implicit Euler time integration method, and the stochastic part was solved by the Wiener process increment with time step, $dt=2.5\times 10^{-3}$.
@@
We quantified tissue circularity by computing the convex hull of the final cell-center positions. From the resulting QHull polygon (Fig. S2), we evaluated its area ($A$) and perimeter ($P$), and defined circularity as
@@
This gives a value of 1 for a perfect circle.
Figure S2: Using convex hull to evaluate the circularity of tissue morphology.
@@
Where $W(x)$ is a Heaviside step function given by $W(x<0)=1$. It quantifies the radial movement of a cell from its initial position. The relaxation time for a tissue is defined by the time it takes for $Q(t)$ to become $\frac{1}{e}^{th}$ of its initial value.
@@
Figure S3: Cumulative count of T1 transitions ocer time. Higher motility promotes T1 transitions.
@@
Where $\Pi_{c}=-\frac{\partial U_{c}}{\partial A_{c}}$ is the hydrostatic pressure and $\boldsymbol{T}_{e}=\frac{\partial U_{c}}{\partial\boldsymbol{l}_{e}}$ is the line tension (or shear stress ) term. This leads to the expression,
@@
$\sum_{v=1}^{N_{c}}\begin{pmatrix}\Delta x^{2}_{v+1}&\Delta x_{v+1}\Delta y_{v+1}\\ \Delta y_{v+1}\Delta x_{v+1}&\Delta y^{2}_{v+1}\end{pmatrix}\frac{1}{\sqrt{\Delta x^{2}_{v+1}+\Delta y^{2}_{v+1}}}$
@@
A stress relaxation experiment involves applying an external deformation to the system and observing how the internal stress dissipates over time. We apply an affine shear strain ${\bf\gamma}(t)=\left(\begin{array}[]{cc}1&\epsilon\\ 0&1\end{array}\right)$ to the system and maintain the shear strain by keeping the boundary of the tissue fixed at a particular level of strain. We examine the relaxation of the shear stress of the bulk tissue, far away from the boundary.
@@
We have applied oscillatory shear strain to the system with varying frequency of the oscillation and measured the stress response of the system. From this, we calculated Storage ($G^{\prime}$) and Loss modulus ($G^{\prime\prime}$), given by the respective in-phase and out-of-phase responses of stress ($\sigma(t)$) to the applied strain ($\epsilon(t)$.
@@
Thus, we have to find out the Fourier coefficients corresponding to the input frequency $\omega_{0}$ from the Fourier series of $\sigma(t)$. One way is to do a Fast Fourier Transform (FFT) numerically and calculate the coefficients. However, this requires extensive sample data and can be prone to error at times. To avoid that, we have used an alternate calculation as follows:
@@
Let’s say $n=n_{0}$ corresponds to the input frequency. So we have to find the coefficients of $\cos(\omega_{n_{0}}t)$ and $\sin(\omega_{n_{0}}t)$ i.e $a_{n_{0}}$ and $b_{n_{0}}$. Now, if we integrate equation S13 to a time $T$ or multiply it with $\cos(\omega_{n_{0}}t)$ or $\sin(\omega_{n_{0}}t)$ and then integrate, we get three equations involving $a_{0}$, $a_{n_{0}}$ and $b_{n_{0}}$ as three unknowns which we can solve as system of linear equations. These equations look like,
@@
In equation S14, the other terms where $n\neq n_{0}$ have been ignored as they would not contribute much. In equation S14, left-hand side integrals are calculated using a numerical integration scheme. Thus, we can calculate $a_{0}$, $a_{n_{0}}$ and $b_{n_{0}}$, and from that we can get $G^{\prime}=\frac{b_{n_{0}}}{\epsilon_{0}}$ and $G^{\prime\prime}=\frac{a_{n_{0}}}{\epsilon_{0}}$. The amplitude of the oscillation was kept sufficiently small ($\epsilon_{0}\ll 1$) to probe a linear response. Simulations with higher motility required a higher ensemble average and low-pass filtering in the resulting stress before calculating the moduli.